At what velocity will you jog to travel 300.0 m after 40.0 seconds?


1
Expert's answer
2021-10-26T09:38:39-0400
"v=\dfrac{s}{t}=\dfrac{300\ m}{40\ s}=7.5\ \dfrac{m}{s}."